OpenStudy (anonymous):

A cylinder is inscribed in a right circular cone of height 2 and radius (at the base) equal to 2. What are the dimensions of such a cylinder which has maximum volume?

OpenStudy (anonymous):

calculus

OpenStudy (anonymous):

This is an answer for a radius of 6... just substitute your 2 for the 6 The radius of the inscribed cylinder at height x is given as (6/2 * (2-x)) so V = pi* (6/2 * (2-x))^2 * x dV/dx = 0 = (2-x)^2*3+3x(2(2-x)(-1)) 3(2-x)^2=6x(2-x) 2-x = 2x 3x = 2 x = 2/3 so the cylinder has a height of 2/3 and radius of (6/2 * (2-x)) = 4

OpenStudy (anonymous):

apparently the radius is wrong :/

OpenStudy (anonymous):

do your own hw, stop asking people to do it for you